Adriana Corporation manufactures football equipment. In planning for next year, the managers want to understand the relation between activity and overhead costs. Discussions with the plant supervisor suggest that overhead seems to vary with labor-hours, machine-hours, or both. The following data were collected from last year's operations:

Month Labor-Hours Machine-Hours Overhead Costs
1 715 1,361 $ 102,723
2 705 1,416 103,803
3 685 1,510 109,937
4 745 1,446 108,257
5 785 1,596 116,210
6 745 1,577 114,410
7 735 1,389 107,052
8 735 1,311 102,017
9 720 1,453 106,391
10 785 1,553 113,049
11 675 1,290 98,865
12 705 1,606 110,873

Required:

a. Use the high-low method to estimate the fixed and variable portions of overhead costs based on machine-hours. (Round "Variable cost" answer to 2 decimal places.)

Variable Cost (per machine hour)
Fixed cost


b. Managers expect the plant to operate at a monthly average of 1,700 machine-hours next year. What are the estimated monthly overhead costs, assuming no inflation? (Round"Variable cost" answer to 2 decimal places.)

Overhead Costs
Answer #1

a) Variable cost per machine hour = Change in Cost/Change in machine hour = (110873-98865)/(1606-1290) = 38 per machine hour

Fixed cost = 110873-(1606*38) = 49845

b) Overhead cost = (1700*38+49845) = 114445